OpenStudy (anonymous):

Please help! Express 1+cos24 as a function of half the given angle. a. 2cos^2 48degrees b. 2cos^2 12degrees c. 4cos^2 48degrees d. 4cos^2 12 degrees

OpenStudy (anonymous):

use this \[\frac{1+\cos 2\theta}{2}=\cos^2 \theta\]

OpenStudy (waleed_imtiaz):

Its Answer: B

OpenStudy (anonymous):

@waleed_imtiaz How did you get B?

OpenStudy (ganpat):

1+cos24.. represent it in (1+cos2θ)/2=cos2θ .. form so multiply n divide by 2.. 2( 1+ cos 2(12)) / 2 ( 1+ cos 2(12)) / 2 = \[\cos ^{2}12.. \] So answer = 2 \[\cos ^{2} 12 degrees..\] option B...

OpenStudy (anonymous):

Thanks guys!

OpenStudy (waleed_imtiaz):

There's a rule , just remember this..... If the left hand side of the equation has the angle 4x(say). then the angle in the right hand side would be half of the left hand side...... i.e., 2x... So from this, U can formulate any angle ..... U can place any angle and get the result in double angle, or half angle identities...... Let's see If I say..... 1+cos48x= ? then according to my above justification, U can write it as =2cos^2(24x) which is the half of 48x... Now any confusion.... U can ask me....... ThanK you.....